What is the magnitude of the momentum of a basketball player weighing 72 kg running towards the back court for a fast break play with a speed of 7m/s?


1
Expert's answer
2021-04-19T17:08:29-0400

Solution. Using the definition of the momentum


"p=mv"

where m=72kg is the mass of the body, v=7m/s is the velocity of the body.

Therefore, the magnitude of the momentum is equal to


"p=72kg\times7\frac{m}{s}=504\frac{kg\times m}{s}"

Answer.

"p=504\frac{kg\times m}{s}"